OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of NOT pulling out a red marble three times if none are replaced? A.35/288 B.21/55 C.6/11 D.210/1728

OpenStudy (anonymous):

I just do not know how to do this.... Having much difficulty

Directrix (directrix):

There are 12 marbles in the container, 9 of which are not red. On the first draw, the probability of drawing a non red = 9/12 On the second draw, the probability of drawing a non red = 8/11 On the third draw, the probability of drawing a non red = 7/10 The probability of not drawing a red marble on the first three draws is: 9/12 * 8/11 * 7/10 = 504/1320 which simplifies to 21/55.